OpenStudy (anonymous):

@JuanitaM i need more help itss like im not sure if im gettin if its cos, sin, or tan hen i check the answer in the back of the book and its wrong

OpenStudy (anonymous):

@dessyk i just answered on the initial post

OpenStudy (anonymous):

sin Angle = opp/hyp cos Angle = adj/hyp tan angle = opp/adj